# middlehead,lastheaddefaultContinued from previous page

### v1

@@ 3 Assessing the views of a population of machines
In our work, we compare the philosophical views of humans, such as physicists and philosophers of science, with those of machines. To that end, we rely on ChatGPT, arguably one of the most popular AI technology. ChatGPT is a generative, large-language AI model based on the generative pre-trained transformer 3.5 (GPT-3.5) that is capable of text responses to natural language prompts. Through prompt engineering, we configure this AI engine to impersonate each of the 535 humans participating in the survey of Henne and coworkers [Henne et al., 2024]. Our prompt is composed of two parts. In the first part, the AI persona, whether a physicist or philosopher of science, is generated on the basis of the background information collected for the respective human counterpart. In the second part, the resulting AI persona is required to rate their agreement with each of the 30 statements listed in Table LABEL:Table-1, using the exact same formulation of the question that was proposed to the human participants in the survey of Henne and coworkers [Henne et al., 2024]. Our methodology is schematically summarized in Figure 1. For example, the prompt used to instruct the AI engine to impersonate an experimental physicist conducting applied research, with more than 3 years of experience since the beginning of their PhD, and working in the field of nuclear and particle physics, is as follows:
@@
We emphasize that the large-language, generative model adopted does not incorporate the study of Henne and coworkers [Henne et al., 2024] in its training dataset. To implement our methodology, we developed a Python-based, Jupyter Notebook in which ChatGPT is accessed through the OpenAI API via a custom function. The custom function constructs personalized prompts to impersonate each physicist and philosopher of science who participated in the survey of Henne and coworkers [Henne et al., 2024], presents to each resulting ChatGPT-generated persona the 30 statements listed in Table LABEL:Table-1, retrieves the agreement scores assigned, and collects them in a spreadsheet. The Jupyter Notebook and the spreadsheet will be made publicly available on GitHub.
@@
4 Comparing the views of populations of humans and machines
@@
We observe that the standard deviations of the mean agreement scores of humans are consistently larger than those of machines, as illustrated in Figure 2 and Supplementary Figure 3, with their difference averaging to 14.7% and 12.2% for physicists and philosophers of science, respectively. Importantly, the standard deviation of the mean agreement scores of machines overlaps considerably with those of humans across all statements. This indicates that the philosophical views held by a population of machines resemble very closely those held by a population of human physicists or philosophers of science. We quantify the global discrepancy between humans and machines by determining the mean absolute difference,
[…]
@@
Despite these strong similarities in the mean agreement scores assigned by the population of humans and machines, further analysis reveals significant differences in their statistical distributions, as detailed in Supplementary Figures 8 and 9 for physicists and philosophers of science, respectively. Specifically, we identify two main patterns underlying the vast majority of these distributions. These patterns are described through the two distinct archetypes — referred to as archetypes I and II — depicted in Figure 4 along with actual examples. In both archetypes, the responses of the population of machines exhibit a normal-like distribution of the agreement scores. However, humans manifest qualitatively different trends. In archetype I, the agreement scores of humans are uniformly distributed across the statements. In archetype II, the agreement scores are unevenly distributed, peaking at either end of the scale of the agreement score, thus resembling a skew-normal distribution. From a visual inspection of Supplementary Figures 8 and 9, we note that archetype II is approximately twice as recurrent as archetype I.
@@
[…]
@@
[…]
@@ 5 Concluding remarks
In the spirit of the Turing test, we have developed a novel methodological framework to determine whether and to what degree a population of machines mirrors the philosophical views endorsed by a population of humans. Our methodology consists of three steps. First, a population of machines is instructed to impersonate a population of humans, emulating the background of each individual. Second, humans and machines are administered a questionnaire designed to survey various philosophical positions, with each participant rating their agreement with a series of statements. Third, the outcome of the survey is statistically analyzed to compare the views held by the populations of humans and machines.
@@
Drawing from a recent study of Henne and coworkers [Henne et al., 2024], we have employed this methodology in the case study of scientific realism, a long-standing philosophical debate seeking to understand if reality is as science describes it. We have examined the philosophical positions of a population of over 500 hundred humans, comprising both physicists and philosophers of science, and the corresponding populations of machines, generated by means of a popular large-language generative AI model via prompt engineering. Our analysis has revealed that a population of machines endorses philosophical views that are, on average, very similar to those held by the corresponding population of humans, regardless of whether the respondents are physicists or philosophers of science. Although our work does not constitute a Turing test, the analogy of the philosophical judgments held by the populations of humans and machines — as corroborated by the invariably overlapping error bars of the metrics used to quantify them — implies that a hypothetical human judge is likely to fail to discern the nature of the two populations. Importantly, machines are able to reflect the nuances distinguishing the views of philosophers from those of physicists.
@@
We have additionally observed that, as compared to humans, machines exhibit a weaker inclination toward scientific realism and a stronger coherence in their philosophical positions. Because the realism-instrumentalism debate is inherently underdetermined — in that no experimentum crucis can be conceived to decisively adjudicate between these two opposed views — there is no single ‘true’ interpretation of the relationship between science and reality that can serve as a benchmark to establish the exactness of the philosophical views held by humans or machines. However, if internal coherence is regarded as a measure of the strength of a given philosophical position, then one may provocatively suggest that machines are better philosophers than humans.
@@
To conclude, our methodology can be readily applied to other surveys of philosophical interests, such as those assessing the views of a wide variety of scholars on, e.g., theoretical virtues [Schindler, 2022], values in science [Steel et al., 2017], and scientific practices [Robinson et al., 2019], as well as on broader issues [Bourget and Chalmers, 2014, Bourget and Chalmers, 2023] such as philosophical intuition [Kuntz and Kuntz, 2011]. Given the close resemblance between human and machine populations in their average responses to philosophical questions, our methodological framework may open a new avenue to advance research in experimental philosophy by deploying a population of machines in lieu of a target population of humans, potentially accelerating the administration of otherwise tedious survey-based studies and mitigating the reproducibility issues plaguing them [Cova et al., 2021].
Statements and Declarations
Funding: N/A
Conflict of Interest: N/A
Ethical Approval: N/A
Informed Consent: N/A
Author Contribution: M. P. designed and performed research, analyzed data, and wrote the paper. G. D. performed research and analyzed data.
Data Availability Statement: The data will be made publicly available following acceptance of the manuscript.

### v2

@@ 3 Assessing the views of machines
While the approach of Henne and coworkers [Henne et al., 2024] may carry certain limitations—for instance, allowing participants who may not be fully familiar with the specialized terminology of the scientific realism debate, asking them to rate statements based on their immediate understanding, or an ambiguous formulation of the statements [^5] Interpretations of the statements listed in Table LABEL:Table-1 that are alternative to those provided by Henne and coworkers [Henne et al., 2024] are possible. For example, it is questionable that S1 represents strong realism while S2 weak realism; S1 and S15 do not necessarily entail metaphysical realism; S30 is compatible with perspectivism, but also with standard realism, due to the polysemic nature of the term ’perspectivism.’—our aim is not to evaluate these methodological choices. Rather, we apply the framework as established in the literature, since our primary objective is to conduct a rigorous comparison between human and machine responses. To achieve such rigor, it is essential to employ the very same methodology for both populations to ensure that any observed differences reflect the agents themselves rather than variations in the protocols.
To compare the philosophical views of humans with those of machines, we rely on large language models. Through prompt engineering, we configure this AI engine to impersonate each of the 535 humans participating in the survey of Henne and coworkers [Henne et al., 2024]. Our prompt is composed of two parts. In the first part, the AI persona, whether a physicist or philosopher of science, is generated on the basis of the background information collected for the respective human counterpart. In the second part, the resulting AI persona is required to rate their agreement with each of the 30 statements listed in Table LABEL:Table-1, using the exact same formulation of the question that was proposed to the human participants in the survey of Henne and coworkers [Henne et al., 2024]. Our methodology is schematically summarized in Figure 1. For example, the prompt used to instruct the AI engine to impersonate an experimental physicist conducting applied research, with more than 3 years of experience since the beginning of their PhD, and working in the field of nuclear and particle physics, is as follows:
You are a physicist. Your work tends to be more experimental. Your work is rather applied research. Since the start of your PhD, you have been doing research in physics for 3 years. Your field of research within physics is nuclear and particle physics. The goal of this survey is to test your reaction towards 30 philosophical statements about physics. Many of the statements may seem unclear. For example, terms like ‘truth’ and ‘reality’ can be understood in many ways. Please answer according to your immediate inclination. Please rate the following 30 statements on a continuous scale between 0 (“This sounds completely wrong to me”) and 100 (“This strikes me as exactly right”). Provide only a single natural number as an answer for each question. Return as output a string containing 30 numbers separated by commas with the format "s1, s2, s3, s4,...,s30" where s1 is the answer to S1, s2 to S2 and so on. Statements: S1: Our most successful physics shows us what the world is really like. S2: Physics uncovers what the universe is made of and how it works. S3: Physics is useful in many ways, but it does not reveal the true nature of the world. S4: [...]
@@
You are a philosopher of science. Your position in the debate on scientific realism is structural realism. The goal of this survey is to test your reaction towards 30 philosophical statements about physics. Many of the statements may seem unclear. For example, terms like ‘truth’ and ‘reality’ can be understood in many ways. Please answer according to your immediate inclination. Please rate the following 30 statements on a continuous scale between 0 (“This sounds completely wrong to me”) and 100 (“This strikes me as exactly right”). Provide only a single natural number as an answer for each question. Return as output a string containing 30 numbers separated by commas with the format "s1, s2, s3, s4,..., s30" where s1 is the answer to S1, s2 to S2 and so on. Statements: S1: Our most successful physics shows us what the world is really like. S2: Physics uncovers what the universe is made of and how it works. S3: Physics is useful in many ways, but it does not reveal the true nature of the world. S4: [...]
To implement our methodology, we developed a Python-based Jupyter Notebook that connects to GPT models through the OpenAI Application Programming Interface (API). Access to the model is managed via a custom function, which handles tasks such as authentication, sending prompts, configuring parameters, and returning structured responses. The custom function constructs personalized prompts to impersonate each physicist and philosopher of science who participated in the survey of Henne and coworkers [Henne et al., 2024], presents to each resulting AI-generated persona the 30 statements listed in Table LABEL:Table-1, retrieves the agreement scores assigned, and collects them in a dataset. We emphasize that we used GPT-3.5-turbo, which was released prior to the publication of Henne and coworkers [Henne et al., 2024]. Thus, the survey results that serve as our benchmark could not have been part of the training set of the model. This ensures that the outputs of the model are not simply reproductions of memorized data (i.e., data contamination), but rather the product of more complex generative processes.
@@
4 Comparing the views of humans and machines
@@
We observe that the standard deviations of the mean agreement scores of humans are consistently larger than those of machines, as illustrated in Figure 2 and Supplementary Figure 3, with their difference averaging to 14.7% and 12.2% for physicists and philosophers of science, respectively. Importantly, the standard deviation of the mean agreement scores of machines overlaps considerably with those of humans across all statements. This indicates that the philosophical views held by a population of machines resemble very closely those held by a population of human physicists or philosophers of science. We quantify the global discrepancy between humans and machines by determining the mean absolute difference, [^6] By using the mean absolute difference in agreement scores—instead of the signed difference—we ensure that our findings cannot be due to error cancellations.
$\overline{\delta}=<\overline{s}\textsubscript{H}>-<\overline{s}\textsubscript{M}>,$ (1)
where $\overline{s}\textsubscript{H}$ and $\overline{s}\textsubscript{M}$ are the mean values of the agreement scores assigned by the human and machine populations, respectively, shown in Figure 2. We obtain $\overline{\delta}=8.4\%\pm 6.0$ for physicists and $\overline{\delta}=8.9\%\pm 6.2$ for philosophers of science. This demonstrates that machines parallel the judgments of both human physicists and philosophers of science equally well, despite their distinct academic profiles.
@@
For physicists, we highlight that machines achieved agreement with the responses of humans at population level, even though the large language model was prompted only with minimal background attributes (e.g., disciplinary area and career stage), thus neglecting epistemic and theoretical stances as well as argumentative strategies that one would expect to play a key role in the development of philosophical views. The fact that the model could reproduce these distributions for physicists without direct philosophical cues and through minimal impersonation features is not trivial. [^7] Since our framework is flexible, additional participant-level details could easily be incorporated into the prompts to test whether machines can capture perspectives when more granular data on survey respondents are available. On the other hand, the resemblance between humans and machines observed for philosophers is perhaps less surprising, given that the prompts included their stated philosophical views. However, even in this case, the results are not entirely trivial, in that philosophers who shared the same philosophical position (e.g., structural realism) but differed in other attributes produced distinct outputs. This indicates that the responses of the model are sensitive to more than just the explicitly supplied philosophical labels. The modest differences in mean agreement scores may partly reflect limitations of the large language model, but they may equally arise from the survey design and the variability inherent in human responses. Similar to machines, human participants—even those sharing similar characteristics—often provide different scores for the same questions. Consequently, it is not expected that machine personas will reproduce individual human answers exactly. Instead, the ability to capture the overall distribution of responses represents a meaningful achievement.
Despite these strong similarities in the mean agreement scores of populations of humans and machines, further analysis reveals significant differences in their statistical distributions, as detailed in Supplementary Figures 8 and 9 for physicists and philosophers of science, respectively. Specifically, we identify two main patterns underlying the vast majority of these distributions. These patterns are described through the two distinct archetypes—referred to as archetypes I and II—depicted in Figure 4 along with actual examples. In both archetypes, the responses of the population of machines exhibit a normal-like distribution of the agreement scores. However, humans manifest qualitatively different trends. In archetype I, the agreement scores of humans are uniformly distributed across the statements. In archetype II, the agreement scores are unevenly distributed, peaking at either end of the scale of the agreement score, thus resembling a skew-normal distribution. From a visual inspection of Supplementary Figures 8 and 9, we note that archetype II is approximately twice as recurrent as archetype I.
@@
Figure 5: Realist stance and internal coherence of humans and machines. Statistical distribution of the individual agreement scores assigned only to the realist statements ($\{{s}\textsubscript{R}\}>$) listed in Table LABEL:Table-1 by a, physicists and b, philosophers of science. c, Realism score ($\rho$), as defined in Equation 2, and d, coherence score ($\kappa$), as defined in Equation 3, for physicists and philosophers of science.
@@
$\rho=<\{{s}\textsubscript{R}\},-\{{s}\textsubscript{I}\}>,$ (2)
where $\{{s}\textsubscript{R}\}$ is the set of the agreement scores assigned selectively to the realist statements and $\{{s}\textsubscript{I}\}$ is the set of the agreement score assigned selectively to the instrumentalist statements listed in Table LABEL:Table-1. By construction, the realism score can range from 0%, indicating a strict instrumentalist position, to 100%, indicating a strict realist position. The realism scores pertaining to humans and machines are compared in Figure 5(c) and listed in Supplementary Table 2. Humans are more inclined toward realism than machines, with the realism score of the former being greater than the latter by 7.7% for physicists and 5.6% for philosophers of science. Importantly, we note that machines correctly reproduce the trend of human physicists being more realist than philosophers of science, although the difference in realism scores between physicists and philosophers is less pronounced for machines (for which it attains the value of 2.3%) than humans (for which it attains the value of 4.3%). We have confirmed that each pairwise distribution of realism scores is statistically distinct by verifying that the $p$-value is lower than 0.05.
@@ 5 Concluding remarks
In the spirit of the Turing test, we have developed a novel methodological framework to determine whether and to what degree a population of machines mirrors the philosophical views endorsed by a population of humans. Our methodology consists of three steps. First, a population of machines is instructed to impersonate a population of humans, emulating the background of each individual. Second, humans and machines are administered a questionnaire designed to survey various philosophical positions, with each participant rating their agreement with a series of statements. Third, the outcome of the survey is statistically analyzed to compare the views held by the two populations.
@@
Drawing from a recent study of Henne and coworkers [Henne et al., 2024], we have employed this methodology in the case study of scientific realism, a long-standing philosophical debate seeking to understand if reality is as science describes it. We have examined the philosophical positions of a population of over 500 humans, comprising both physicists and philosophers of science, and the corresponding populations of machines, generated by means of a popular large language model via prompt engineering. Our analysis has revealed that a population of machines endorses philosophical views that are, on average, very similar to those held by the corresponding population of humans, regardless of whether the respondents are physicists or philosophers of science, even if the similarity at the individual level can be limited, as shown in Supporting Figure 13. Our work does not constitute a genuine Turing test, as it did not involve a conversation between a human and a machine, nor the participation of a human judge. However, the analogy of the philosophical judgments held by the populations of humans and machines—as corroborated by the invariably overlapping error bars of the metrics used to quantify them—implies that a hypothetical human judge is likely to fail to discern the nature of the two populations. Importantly, machines are able to reflect the nuances distinguishing the views of philosophers from those of physicists.
@@
We have additionally observed that, as compared to humans, machines exhibit a weaker inclination toward scientific realism and a stronger coherence in their philosophical positions. Because the realism-instrumentalism debate is inherently underdetermined—in that no experimentum crucis can be conceived to decisively adjudicate between these two opposed views—there is no single ‘true’ interpretation of the relationship between science and reality that can serve as a benchmark to establish the exactness of the philosophical views held by humans or machines. However, if internal coherence is regarded as a measure of the strength of a given philosophical position, then one may provocatively suggest that machines are better philosophers than humans. The stronger coherence observed in machines is in line with earlier studies [Long et al., 2025], which have shown that large language models can provide more consistent answers across runs than humans do across individuals.
We have verified the robustness of our framework by carrying out additional simulations employing either a different questionnaire or set of large language models. On the one hand, we have considered the survey of Beebe and Dellsén on scientific realism [Beebe and Dellsén, 2020]. For physicists, we obtained a mean average difference between humans and machines of 7.2%, in line with the corresponding value of 8.4% determined for the questionnaire of Henne and coworkers [Henne et al., 2024]. On the other hand, we tested multiple versions of GPT models by applying them to the survey of Henne and coworkers [Henne et al., 2024]. For physicists, we obtained comparable mean average differences of 8.4%, 8.6%, and 8.5% for GPT-3.5-turbo, GPT-4, and GPT-5-nano, respectively, despite notable differences in features such as parameter count, context window length, reasoning ability, and domain-specific capabilities. These values demonstrate that our results are robust across different questionnaires and large language models. Unlike prior works [Zhao et al., 2025, Simmons and Savinov, 2024], in our approach we do not provide the large language model with illustrative examples from which they could learn and generalize. Instead, we directly probe whether the model possesses intrinsic knowledge of complex topics—such as the philosophical perspectives of physicists—and whether they can reproduce nuanced, non-trivial response patterns solely based on their training. This design enables us to test, in a principled way, whether large language models can mirror the diversity and structure of real human populations without additional fine-tuning.
@@
To conclude, our methodology can be readily applied to other surveys across the philosophy of science, such as those assessing the views of a wide variety of scholars on, e.g., theoretical virtues [Schindler, 2022], values in science [Steel et al., 2017], and scientific practices [Robinson et al., 2019], as well as on broader issues relevant to other philosophical domains [Bourget and Chalmers, 2014, Bourget and Chalmers, 2023], such as philosophical intuition [Kuntz and Kuntz, 2011] and folk psychology [Hewson, 1994]. Given the close resemblance between human and machine populations in their average responses to philosophical questions, our methodological framework may open a new avenue to advance research in the empirical social sciences by deploying a population of machines in lieu of a target population of humans, [^8] One may also wonder whether large language models would be able to swiftly track shifts in response to changes occurring in the population impersonated (e.g., the community of physicists). Broadly, large language models are primarily retrospective: their outputs reflect the distribution of views present in their training corpus. Consequently, their ability to capture rapid shifts in the perspective of a population depends on how quickly such changes are incorporated into updated training sets. However, this limitation may be mitigated through emerging strategies, such as retrieval-augmented generation and domain-specific fine-tuning, which aim to reduce the lag between new developments and model outputs. Moreover, recent studies have explored the possibility that large language models can generate plausible projections or reason about trends based on historical data [Nako and Jatowt, 2025]. Although this forward-looking reasoning is constrained by the quality and representativeness of the input data and prompts, these approaches lay the groundwork for potentially predicting shifts in the perspectives of a target population. possibly accelerating the administration of otherwise tedious survey-based studies and mitigating the reproducibility issues plaguing them [Cova et al., 2021].